# PersonaFlow: Boosting Research Ideation with LLM-Simulated Expert Personas

### v1

@@ 4.2.2. Understanding User Perception: Post-session survey for each condition and exit interview
Post-Session Survey Participants provided an overall evaluation of the co-creation experience and outcomes at the end of each task through a survey. The survey includes a set of 5-point Likert scale questions intended to assess participants’ experience in terms of ideation outcome quality, ideation experience, and system usability. The complete list of survey questions can be found in Appendix 6.
@@
Figure 5. An example screenshot showcasing the two embedded ratings (the critique and RQ node, from left to right) used to collect users’ perceptions on the fly. Users were instructed to rate at least one node out of any newly generated batch of critique or RQ nodes.
Figure 6. Average of users’ 5-point Likert scale ratings towards the generated critiques and RQs: comparison between different dimensions of ideation outcomes between single- and multi-persona conditions. The ratings are collected on the fly through rating forms embedded in each node during the use of the system, as detailed in section 4.2.1.
@@ 5. Findings
Figure 7. Post-session survey results comparison between single and multiple persona conditions. The x-axis denotes the percentage of Likert scale ratings given by participants by value. A t-test showed that users’ overall perceptions of critique helpfulness significantly improved under the multi-persona condition when compared to single-persona ($t=2.36,p=0.033^{*}$).
5.1. The Impact of Multi-Personas on the Perception of Generated Feedback and User Experience (RQ1)
@@
5.1.1. Compared to single-persona, the multi-persona condition significantly improved users’ ratings of the generated critiques.
@@
Table 1. Number of edits user performed over expert personas and examples by different traits.
5.1.2. Positive correlation between the number of personas and the perceived creativity and specificity of the generated RQs.
We also found a positive correlation between the number of personas used during ideation and the perceived creativity ($r=0.37,p=0.041^{*}$) and specificity ($r=0.29,p=0.014^{*}$) of research questions, indicating a positive impact of LLM-simulated personas on users’ perception of ideation outcomes. Qualitative findings revealed that personas introduced new factors or concepts that users had not previously considered. One participant noted, “… Personas added important factors such as behavior psychology, into the final research questions …” (P1). In other examples, participants emphasized that the personas’ critiques introduced helpful new concepts for understanding a new domain, such as “generative item retrieval”, “ranking”, and “user preference modeling” (P3), or “the impact of cognitive load” during the exploration of the topic “LLM assistant on AR smart glasses” (P5). Additionally, personas influenced the direction of exploration regarding topics. One participant stated, “… help drive more directions like ’AI ethics,’ the exploration direction is changed and shifted …”(P9). The participant specifically noted that some literature might not be directly relevant to the research question, but it provides a broader view outside their own knowledge. This information was drawn from the expertise of the persona, the views and perspectives of the community, and the background of the persona which is helpful in generating this type of literature.
5.1.3. Compared to single-persona, multi-personas did not change users’ overall perceptions of the system’s usability.
We collected user’s post-session survey ratings to understand their perception of the overall ideation process. We found a significant difference ($t=2.36,p=0.033^{*}$) between user-perceived critique helpfulness between using a single persona ($M=3.78,SD=0.83$) and multiple personas ($M=4.56,SD=0.53$), as shown in Figure 7. Usability-related items were identified to be of no significant difference between single and multi-persona, as using the system under the multi-persona condition did not incur significant changes in perception of system functionality and usage experience. However, we also observed a higher level of user-perceived reliance on the system when multiple personas were involved during ideation, although not significant.
We analyzed users’ behavioral and think-aloud data to understand why the use of multiple personas was perceived as beneficial by participants. We observed that users (N=3) who were highly familiar with the chosen topic reported that the system helped them articulate their ideas, which might have previously been implicit. P10 identified a research question that they have been considering and discussing with other researchers, but the system provided helpful information for them to elicit the method to be used and additional factors to consider during evaluation. The participant noted that “… it definitely gave new information (which) really helped me solidify which approach to use … ,” leading to a clearer research idea and a more structured plan for their study. We also observed that another user (P17) manually added a new persona, “Basic Science Educator Mentor,” after encountering the “Clinical Training Mentor” persona. This decision was driven by their understanding of the medical training process, which involves two distinct stages: foundational science comprehension and practical skill application, as stated by the participant. P17 expressed the need for insights from both the “Clinical Skill Expert” and the “Assessment Specialist” to gain a holistic perspective that aligns with their exploration scope. This approach allowed the participant to consider different viewpoints, especially recognizing that users’ questions often emerge from their specific domain expertise. Moreover, another participant (P8) emphasized the importance of incorporating multiple personas in their study, stating that “… (the expertise of) the other two personas are needed for the study to proceed … .” This reflects how the user rationalizes the selection of personas based on the need for a combination of different skill sets in terms of the feasibility of research project. Participants also found the breakdown of expertise into subsequent personas particularly helpful when conducting divergent exploration. P19 mentioned that “It is too much information if all expertise is within one persona … Breakdown of personas is more helpful for dig deep, and choose the direction to go forward … .”
5.2. Understanding the Behavior of Users Interacting with the System (RQ2)
5.2.1. Patterns in Users’ Customization of LLM-Generated Personas
When examining user interaction with the system (RQ2), we focus on how users engaged with LLM-generated personas by customizing and editing the persona profiles based on that initially suggested by the system.
Dependent Variables
*: p$<$0.05, **: p$<$0.01, ***: p$<$0.001

### v2

@@ 4.2.2. Understanding user perception: post-session survey for each condition and exit interview
Post-Session Survey Participants provided their overall evaluation of the co-creation experience and outcomes through a survey. The survey includes a set of 5-point Likert scale questions intended to assess participants’ experience in terms of ideation outcome quality, ideation experience, and system usability. The complete list of survey questions can be found in Appendix 6.
@@
We also further analyzed their behavior and think-aloud data and presented case studies demonstrating how they interacted with the system and how the system promoted their critical thinking activities. We analyzed users’ cognitive activities through the lens of critical thinking (Gopinathan et al., 2017; Facione, 1990; Facione et al., 2011), as critical thinking is crucial to the process of interdisciplinary research ideation (Newton, 2010). We refer to the following critical thinking framework (Facione et al., 2011) during analysis:
  1. (1)
    Interpretation: To comprehend and express the meaning or significance of a wide variety of experiences, situations, data, events, judgments, conventions, beliefs, rules, procedures, or criteria.
  2. (2)
    Analysis: To identify the intended and actual inferential relationships between different arguments and the implicit assumptions in the reasoning.
  3. (3)
    Evaluation: To assess the credibility of statements or other representations that are accounts or descriptions, and to assess the logical strength of the actual or intended inferential relationships among statements, descriptions, questions, or other forms of representation.
  4. (4)
    Inference: To identify and secure elements needed to draw reasonable conclusions, to form conjectures and hypotheses, to consider relevant information, and to reduce the consequences flowing from evidence.
  5. (5)
    Explanation: To state and justify one’s opinions and arguments based on reasons and awareness of possible counterarguments, and to present one’s reasoning in the form of cogent arguments.
  6. (6)
    Self-regulation: To self-consciously monitor one’s cognitive activities, the elements used in those activities, and the results educed, particularly by applying skills in analysis, and evaluation to one’s own inferential judgments with a view toward questioning, confirming, validating, or correcting either one’s reasoning or one’s results.
@@ 5. Findings
5.1. Users’ Perceptions of Experiences and Outcomes in Research Ideation (RQ1)
@@
5.1.1. Multi-perspectives improved users’ perceived Relevance and Helpfulness of the generated critiques.
@@
Figure 6. The left figure is an example screenshot showcasing the embedded rating surveys (the critique node) used to collect users’ perceptions on the fly. Users were instructed to rate at least one node out of any newly generated batch of critique or RQ nodes. The right plot presents the average of users’ 5-point Likert scale ratings towards the generated critiques and RQs: a comparison between different dimensions of ideation outcomes between the earlier stage, where the user interacted with one single expert, and the latter stage, where multiple expert perspectives were introduced. The ratings are collected on the fly through rating forms embedded in each node during the use of the system, as detailed in section 4.2.1.
[…]
Two components: Critique Node Ratings (left) and Research Question (RQ) Node Ratings (right). Each node contains sliders to rate specific statements on a scale from ‘Strongly Disagree’ to ‘Strongly Agree’. The Critique Node allows users to rate aspects such as relevance, helpfulness, information richness, and whether the critique provides new insights. The RQ Node is used to evaluate research questions based on relevance to the research topic, provision of new insights, feasibility for further research, and clarity. These ratings and sliders are magnified. charts of users’ 5-point Likert scale ratings for Critique Ratings and RQ Ratings across two conditions: Single Persona and Multiple Personas. In the Critique Ratings chart, four categories are displayed: relevance, helpfulness, informativeness, and insightfulness. For relevance, the bar for the Single Persona condition is slightly below 4, while the Multiple Personas bar is closer to 4.5, with a p-value of .021 noted above. In the helpfulness category, the Single Persona condition is rated just above 3.5, while Multiple Personas reaches around 4, with a p-value of .016 indicated above the bars. Informativeness and insightfulness show similar ratings for both conditions, around 3.5, with the Multiple Personas bars being slightly higher. In the RQ Ratings chart, the four categories include relevance, creativity, feasibility, and specificity. Both conditions show nearly identical ratings for relevance, with both bars around 4.5. In creativity, feasibility, and specificity, ratings are similar between Single Persona and Multiple Personas, with both conditions consistently around 4.
@@
5.1.2. Positive correlation between the diversity of perspectives and the perceived creativity and specificity of the generated RQs.
We also found a positive correlation between the number of expert perspectives the user engaged with during ideation and the perceived creativity ($r=0.37,p=0.041^{*}$) and specificity ($r=0.29,p=0.014^{*}$) of research questions, indicating a positive impact of LLM-simulated personas on users’ perception of ideation outcomes. Qualitative findings revealed that expert perspectives introduced new factors or concepts that users had not previously considered. One participant noted, “… Personas added important factors such as behavior psychology, into the final research questions …” (P1). In other examples, participants emphasized that the personas’ critiques introduced helpful new concepts for understanding a new domain, such as “generative item retrieval”, “ranking”, and “user preference modeling” (P3), or “the impact of cognitive load” during the exploration of the topic “LLM assistant on AR smart glasses” (P5). Additionally, personas influenced the direction of exploration regarding topics. One participant stated, “… help drive more directions like ’AI ethics,’ the exploration direction is changed and shifted …”(P9). This information was drawn from the expertise of the expert, the views and perspectives of the community, and the background of the expert which is helpful in generating this type of literature.
Figure 7. Post-session survey results comparisons between single and multiple persona conditions. The x-axis denotes the percentage of Likert scale ratings given by participants by value. The post-survey questions can be found in table 6. Ratings 1-5 indicate strong disagreement to strong agreement, respectively. A t-test showed that users’ overall perceptions of critique helpfulness significantly improved under the multi-persona condition when compared to single-persona ($t=2.36,p=0.033^{*}$).
[…]
Stacked bar charts showing post-session Likert scale ratings for Experience Measures, Usability Measures, and Outcome Measures across two conditions: Single Persona and Multiple Personas. The X-axis represents the percentage of ratings, while each bar is divided into five rating categories (Rating 1 through Rating 5). In the Experience Measures chart, categories include Informed New Perspectives, Changed Way of Thinking, Trust in System, Reliance on System, Perceived Control, and Recall of Ideas. The Multi Persona condition shows higher percentages of Rating 5 in categories like Informed New Perspectives and Recall of Ideas, while the Single Persona condition exhibits a wider spread across lower ratings. The Usability Measures chart includes Willingness to Use, Ease of Use, Confidence in Use, and Mental Demand. Multi Persona generally shows higher percentages of Rating 5 in Willingness to Use and Ease of Use, while Mental Demand has more weight in lower ratings (Rating 1 and Rating 2) for both conditions, indicating lower mental demand. The Outcome Measures chart shows Persona Helpfulness, Critique Helpfulness, RQ Creativity, Research Outlines Quality, and Outlines Usefulness. Critique Helpfulness shows a significant difference, with Multi Persona having a higher percentage of Rating 5, indicated by a p-value of .033, while the Single Persona condition has a wider distribution across lower ratings. Other categories show more balanced distributions between conditions.
5.1.3. Engaging with multiple expert perspectives did not increase users’ perceived cognitive load.
We collected users’ post-session survey ratings to understand their perception of the overall ideation process. We found a significant difference ($t=2.36,p=0.033^{*}$) between user-perceived critique helpfulness before ($M=3.78,SD=0.83$) and after ($M=4.56,SD=0.53$) engaging with multiple expert perspectives, as shown in Figure 7. Usability-related items were identified to be of no significant difference before and after users engaged with multiple perspectives, as using the system after multiple perspectives were introduced did not incur significant changes in users’ perception of system functionality and usage experience. However, we also observed a higher level of user-perceived reliance on the system when multiple expert perspectives were involved during ideation, although not significant.
We analyzed users’ behavioral and think-aloud data to understand why the use of multiple expert perspectives was perceived as beneficial by participants. We observed that users (N=3) who were highly familiar with the chosen topic reported that the system helped them articulate their ideas, which might have previously been implicit. P10 identified a research question that they have been considering and discussing with other researchers, but the system provided helpful information for them to elicit the method to be used and additional factors to consider during evaluation. The participant noted that “… it definitely gave new information (which) really helped me solidify which approach to use … ,” leading to a clearer research idea and a more structured plan for their study. We also observed that another user (P17) manually added a new expert, “Basic Science Educator Mentor,” after encountering the “Clinical Training Mentor” persona. This decision was driven by their understanding of the medical training process, which involves two distinct stages: foundational science comprehension and practical skill application, as stated by the participant. P17 expressed the need for insights from both the “Clinical Skill Expert” and the “Assessment Specialist” to gain a holistic perspective that aligns with their exploration scope. This approach allowed the participant to consider different viewpoints, especially recognizing that users’ questions often emerge from their specific domain expertise. Moreover, another participant (P8) emphasized the importance of incorporating multiple expert perspectives in their study, stating that “… (the expertise of) the other two personas are needed for the study to proceed … .” This reflects how the user rationalizes the selection of personas based on the need for a combination of different skill sets in terms of the feasibility of research project. Participants also found the breakdown of expertise into subsequent expert perspectives particularly helpful when conducting divergent exploration. P19 mentioned that “It is too much information if all expertise is within one persona … Breakdown of personas is more helpful for dig deep, and choose the direction to go forward …”
5.2. Case Studies: PersonaFlow Design Supporting Diverse Critical Thinking Activities (RQ2)
Based on the findings of RQ1, we identify potential shifts in users’ ways of thinking before and after their engagement with multiple expert perspectives. In order to further understand how users utilized the systems and the implications of providing diverse perspectives, we first analyzed their critical thinking activities reflected through their think-aloud data during the exploration. Additionally, we examined the association between user behaviors during system use and their overall perception of the ideation process by performing a linear regression analysis using the system logs and correlating the quantitative findings with users’ think-aloud data.
After the task-based study session, we invited participants to an optional 30-minute free exploration session where they freely used the system’s features without guidance or constraints. Researchers observed their usage of different nodes. This section analyzes their behavior and think-aloud data (N=10) by combining data collected from the exploratory study with the open-ended exploration session, and presents cases demonstrating how they collaborated with AI to co-create research ideas and questions. We analyzed how the system promoted users’ cognitive thinking process, more specifically critical thinking skills, as defined in section 4.2.3. We found that most users engaged in critical thinking activities including interpretation, analysis and evaluation. In this section, we highlight case studies demonstrating how participants moved beyond these critical thinking processes and utilized the system to support unique critical thinking activities.
Figure 8. P20’s thought processes: The user started by exploring different personas and RQ nodes broadly, and then decided to dive deep into interacting with one “remote” expert profile (“Spatial Gene Profiler”) with the topic “single-cell RNA sequencing and spatial transcriptomics.” Though the expert was not immediately related to the P20’s research, the recommended literature was surprisingly relevant to P20’s work and exactly what they read recently. This “synergy” quickly elicited P20’s interest in a deeper exploration of this expert. A larger version of the workflow screenshot can be found in section A.4.
[…]
The diagram is divided into four panels, each representing different stages of P30’s interaction with persona and RQ nodes over multiple iterations. The panels are labeled as Iteration #1-3, Iteration #4, Iteration #5, and Iteration #6, showing a progressive flow of exploration across different topics and personas. In the first panel (Iteration #1-3), the user explores various persona and RQ nodes broadly without focusing on a specific topic. By Iteration #4, the user engages with the more focused ‘Spatial Gene Profiler’ persona, initially unrelated to P30’s work, but offering relevant literature aligned with their interests. In Iteration #5 and #6, the user delves deeper into this topic, finding the suggested research highly relevant for further interaction. The flow of nodes, marked by curved arrows across the four panels, demonstrate the deepening focus across iterations.
5.2.1. Typical critical thinking activities (Interpretation, Evaluation, and Inference) applied for ideating RQs
Figure 8 shows an example of participant P20 generated a flow with a depth of six iterations. The participant first examined all expert perspectives generated using P20’s initial research question. At the beginning of using the system, P20 interpreted their own context and needs by reviewing all expert perspectives to see which persona node aligned with P20’s domain in bioinformatics and cancer research. Among the initial expert perspectives, “Bio Data Interpreter” quickly drew the participant’s attention because its literature recommendations matched some publications P20 had already read. This led P20 to evaluate the system’s credibility, “Some of the papers are exactly the same paper I just read… that’s pretty interesting),” strengthening the participant’s trust in the tool. In the first three rounds, P20 mostly analyzed expert perspectives that were closely aligned with their existing research focus. However, P20 then became curious about a persona labeled “Spatial Gene Profiler.” Although this expert perspective was slightly outside the participant’s typical scope, P20 recognized unexpected overlaps after examining its suggested articles. “Some of the papers I didn’t expect to be relevant actually linked back to my research.” Based on the new information, P20 inferred a possible link to their domain, triggering a deeper dive into “single-cell RNA sequencing and spatial transcriptomic,” which is a more specific topic than the participant’s original broad question on “spatial distribution of RNA.”
Figure 9. P5’s thought-process (displaying 15 out of 63 nodes). The participant revised their initial research question after reviewing the generated expert profiles and literature retrieved based on the experts’ perspectives.
[…]
The diagram demonstrates Participant T5’s partial block flow. It includes a Research Question node where the participant initially asked, ‘How to develop a wearable AI LLM (Large Language Model) assistant on AR smart glasses to help UX research conduct and record field study or shadowing?’. This node is circled and magnified, and a note indicates that the user edited this original research question after reviewing the generated personas. The revides research question indicates ‘How to develop a wearable AI LLM(Large language model) assistant on AR smart glasses to help UX research conduct & record field study or shadowing (experimenter observe the participants in shadow) study’. The right panel displays various personas generated from the initial research question. Multiple Literature nodes are displayed which are linked to the previous Persona Nodes. Additional Critique nodes are linked to the previous Literature Node. This flow illustrates how the participant iterated on their research question after exploring the personas and literature nodes, refining the question to focus more on the development of a wearable AI LLM for AR smart glasses in UX research.
5.2.2. Self-regulation activities triggered to reflect on the initial RQ
Figure 9 depicted an example where the participant went back and re-framed their original RQs after evaluating the expert profiles and critiques, to better align the RQs more closely with their evolving exploration interests. This was a common observation among the participants. Participant P5 first generated three expert profiles from their initial RQ node. After reviewing the literature nodes, the user found them-self interested in none of the suggested literature. The participant reflected “… in our language or in our lab, we typically talk about AI assistant, but here [it] doesn’t show any [key]words about AI … I feel there’s a misalignment between human and the AI agent … So I need to make my words more specific,” indicating self-regulation by re-framing and clarifying their initial idea after reading system-generated output. As a result, the participant revised the original research question from “How to develop a wearable assistant on AR smart glasses to help UX research conduct & record field study or shadowing study?” to “How to develop a wearable AI LLM (Large language model) assistant on AR smart glasses to help UX research conduct & record field study or shadowing (experimenter observe the participants in shadow) study?” The revision led to three additional expert perspectives, two of which the participant found helpful, prompting further exploration through follow-up literature suggestions. Specifically, the participant found the newly generated persona “UI Research Expert,” with experience in “field studies and user shadowing,” particularly helpful in providing relevant literature. Additionally, the persona “AI LLM Expert,” specializing in “designing user-friendly interfaces for AR devices,” closely aligned with their exploration interests. We also observed another instance of self-regulation activity, where the user commented “… it’s suggesting what I was thinking. I trust it because it aligns with what I know …” by reflecting on the critique suggestions provided by one of the experts.
5.3. Relationships between PersonaFlow Interactions and User Perceptions (RQ3)
As shown in section 5.2, we observed participants engaged in critical thinking activities when interacting with expert perspectives suggested by the system. In order to uncover deeper insights into users’ usage patterns and perceptions of the system, we further analyzed their behavior data through system logs.
5.3.1. Feature usage impacted users’ perceived experience, outcome and usability.
In this section, we explored how users’ behavior, such as modifications and adaptations made by users to the expert profiles, influence the outcomes of the research questions. We analyzed the association between users’ behavior factors and their perception of the ideation process measured through ratings collected through the post-session survey. A set of linear regression analyses was conducted with post-session survey ratings as dependent variables and users’ behavioral factors (including the number of edits over RQ/critique/persona nodes and the percentage of nodes used by the user for subsequent generation) as predictors (independent variables). We measured the user’s percentages of nodes used by computing the percentage of nodes used for generating subsequent nodes out of all nodes generated. The regression analysis results are shown in Table 1.
Dependent Variables
*: p$<$0.05, **: p$<$0.01, ***: p$<$0.001
[…]
The Table presents regression results for overall survey items based on independent variables such as RQ node edit count, Critique node edit count, Persona node edit count, percentage of nodes used, and total node count. These findings indicate how different edits and node usage impact survey item responses. Significant findings include a positive impact of the percentage of nodes used (β = 2.472, p ¡ 0.01) and total node count (β = 0.041, p ¡ 0.001) on Critique Helpfulness, while RQ Creativity shows a negative association with RQ edits (β = -0.376, p ¡ 0.05). Research Outline Quality is positively associated with the percentage of nodes used (β = 3.015, p ¡ 0.01). Reliance on System has a negative relationship with RQ edits (β = -0.396, p ¡ 0.01). Both Perceived Control and Recall of Ideas show positive relationships with persona edits (β = 0.149, p ¡ 0.01) and (β = 0.185, p ¡ 0.01), respectively, while Mental Demand is negatively associated with total node count (β = -0.034, p ¡ 0.01) and positively with persona edits (β = -0.217, p ¡ 0.05).
5.3.2. Editing the RQ Nodes lowered users’ perceived reliance on the system
Our findings showed that users’ number of edits on RQ nodes was negatively associated with the overall perceived creativity of RQ nodes ($t(36)=-2.060,p=0.039^{*}$), the perceived extent to which the system changed their way of thinking ($t(36)=-3.548,p<0.001^{***}$) and reliance on the system ($t(36)=-3.164,p=0.002^{**}$). The number of users’ edits over RQ nodes was negatively associated with the creativity of RQs, their perceived extent to the system changed their ways of thinking, as they identified the need to modify RQ nodes when they felt that the generated research questions were inadequate or misaligned with their ideation goals. Thus by allowing users to make edits over the RQ nodes, the system provided users with greater control over the content, which consequently led to a lower perceived reliance on the system.
5.3.3. A higher % of nodes used associated with the improved quality of critiques and outlines.
We found significantly positive associations between users’ percentages of nodes used (of Nodes Used) and their ratings given towards the helpfulness of generated critiques (Critique Helpfulness) during the post-session survey ($t(36)=3.070,p=0.002^{**}$). A similar positive association also exists between the total number of nodes created by each user (Total Node Count) and their critique helpfulness (Critique Helpfulness) survey ratings ($t(36)=3.639,p<0.001^{***}$). The results indicate that as users explored more nodes, they received higher-quality critiques from personas. As mentioned by the participants, new insights were discovered through reading critiques during the exploration. P13 mentioned that, “the system provided a new approach for exploration that I hadn’t initially considered,” emphasizing the importance of diverse exploration in enhancing the quality of user-perceived critiques.
5.3.4. More edits on expert profiles improve users’ perceived recall of the generated ideas and perceived control during use.
More edits made on expert profiles (# of edits on Persona nodes) led to users’ enhanced recall (Recall of Ideas) of the generated ideas ($t(36)=2.830,p=0.005^{**}$), As shown in Table 1. Users’ perceived control of the system (Perceived Control) was also improved through more edits over expert profiles (# of edits on Persona nodes) during the exploration ($t(36)=2.617,p=0.009^{**}$). The perceived relevance and helpfulness of critiques did not show a significant correlation with users’ expert profile edits. While we cannot state that expert profile edits enhance ideation outcomes, our findings suggest that utilizing a higher percentage of nodes contributes to fostering a stronger sense of recall over the ideas co-created between the user and the system and allows users to have a stronger sense of agency in shaping the generated ideas. Some participants also mentioned the mind-map-based design specifically helped them identify and recall system-generated research ideas, “… I personally also create workflows to have a clear direction (during research) or just brainstorming … it helped me recall and identify the ideas I had during the process … ” as noted by P9.
5.3.5. Users’ customization of LLM-generated expert profiles.
To provide a more in-depth understanding of participants’ customization interactions over expert profiles, we qualitatively examined system logs showing users’ edits over expert profiles. We focus on understanding how users engaged with LLM-generated personas by customizing and editing the persona profiles based on what was initially suggested by the system. An overview of all edits made by participants are presented in table 2.
@@ 5.3.2. Editing the RQ Nodes lowered users’ perceived reliance on the system
Table 2. Number of edits user performed over expert personas and examples by different traits, using the interface showcases in fig. 2.
[…]
The table summarizes the number of edits made by users to expert personas, organized into two main categories: Expertise and Roles. The Expertise category includes seven traits: Experience, Domain, Skills, Method, Education, and Knowledge. Each trait shows an example of a user modification, with the original text struck through in grey and the new text added in green. For example, under Experience, “12 years in AI research and development” is replaced with “AI expert in large language models, with HCI background.” The total number of edits per trait is listed, with 14 edits for Experience, 12 for Domain, 6 for Skills, 2 for Method, 2 for Education, and 4 for Knowledge. The Roles category includes three traits: Functional Role, Social Identity, and Social Status. Examples of modifications are shown, such as changing ”Assistive Technology Developer” to ”Researcher” under Social Identity, with a total of 19 edits for Functional Role, 11 for Social Identity, and 3 for Social Status.